# BNSR-Invariants of Surface Houghton Groups

### v1

@@ 5 Subgroups of maximal finiteness length
We use the notation $fl(G)=n$ to mean that $G$ is a group of type $F_{n}$ but not of type $F_{n+1}$. We know that if $G<H$ is finite index, then $fl(G)=fl(H)$. But what of the converse, that is, when does $fl(G)=fl(H)$ imply that $G$ is finite index in $H$? For Houghton groups, and for the (pure) surface Houghton groups, the answer is positive for sufficiently large subgroups, in particular for coabelian subgroups. We denote by $G^{\prime}$ the commutator subgroup of a group $G$. Recall that the commutators of the Houghton group $H_{n}$ and the pure surface Houghton group $P\mathcal{H}_{n}$ are the finitely supported and compactly supported elements, respectively. (Really, we could say compactly supported in both cases, using the discrete topology for the $\mathbb{N}$-rays on which the Houghton group acts.) For Theorem 5.1 and Proposition 5.2, let $H$ be either $H_{n}$ or $P\mathcal{H}_{n}$.
@@
Let $G<H$ be a subgroup intersecting the commutator in a finite index subgroup. If $fl(G)=n-1$, then $G$ is finite index in $H$.
Note that it suffices to show for groups containing the commutator. We begin by showing that when $G$ contains the commutator, the image of $G$ in the abelianization is a maximal rank sublattice, which will then imply finite index.
@@
Let $G<H$ be as above, i.e. $fl(G)=n-1$, and $G$ contains the commutator $G^{\prime}$. Write $F$ for the abelianization map to $\mathbb{Z}^{n-1}$. Then $F(G)$ is finite index in $\mathbb{Z}^{n-1}$.
@@
We now prove Theorem 5.1, using the characterization of the commutator as the elements of compact support. Note that for the Houghton group, the ends correspond to the rays in the obvious way.
@@
Suppose first that $G$ contains the commutator. Consider the map of coset spaces $\Psi:H/G\rightarrow\mathbb{Z}^{n-1}/(\textbf{c}\mathbb{Z}^{n-1})$ given by $hG\mapsto F(h)\textbf{c}\mathbb{Z}^{n-1}$. To see that this is well-defined, suppose $h_{1}G=h_{2}G$: then $h_{1}=h_{2}g$ for some $g\in G$, and $F(h_{1})=F(h_{2})+F(g)$. As $F(g)\in\textbf{c}\mathbb{Z}^{n-1}$, we see that $\Psi(h_{1}G)=\Psi(h_{2}G)$. We now wish to show that $\Psi$ is injective. Suppose that $\Psi(h_{1}G)=\Psi(h_{2}G)$. Then $F(h_{1}h_{2}^{-1})=F(h_{1})-F(h_{2})=F(g)$ for some $g\in G$. So our question becomes: does $F(h_{1}h_{2}^{-1})\in F(G)$ imply $h_{1}h_{2}^{-1}\in G$? The element $h_{1}h_{2}^{-1}g^{-1}$ will have no net translation in any end, and hence is compactly supported, and therefore is in $G$.
As there are clearly subgroups of finite index in both cases (take the pre-images of a finite index subgroup of $\mathbb{Z}^{n-1}$, there is one loose end remaining: the infinite index case. Specifically, do there exist subgroups $G<H_{n}$ (or $G<P\mathcal{H}_{n}$) whose intersection with the commutator have infinite index, and such that $fl(G)=n-1$. First, we see that $G$ is necessarily infinite index in $H_{n}$ (or $P\mathcal{H}_{n}$), so all that remains is to check whether this case actually occurs. We thank Noel Brady for the proofs of the following lemmas.
@@
Suppose $[H:G]=m<\infty$. Write $H=Gh_{1}\cup\cdots\cup Gh_{m}$, as a disjoint union. For each index $i$ such that $K\cap Gh_{i}\neq\emptyset$, there is some $g_{i}\in G$ and $k_{i}\in K$ such that $k_{i}=h_{i}g_{i}$. Then we can write
@@
[…]
For the converse, consider a decomposition $H=Gh_{1}\cup Gh_{2}\cup\cdots$. Intersecting this with $K_{i}$, we see as before that the index $[K_{i}:G\cap K_{i}]$ is the numebr of indices $j$ such that $Gh_{j}\cap K_{i}$ is non-empty. As every $h_{j}$ must be in some $K_{i}$, we see that if $[H:G]$ is infinite, then the sequence $([K_{i}:G\cap K_{i}])$ goes to infinity. ∎
We return to our question on the existence of infinite index subgroups of Houghton groups with finiteness length $n-1$, We must now split off the ordinary Houghton group from the surface Houghton group. For $H_{n}$, the answer to our existence question is easily yes: there are even copies of $H_{n}$ itself as infinite index subgroups of $H_{n}$! Consider the stabilizer of a single point in $[n]\times\mathbb{N}$: ignoring the fixed point, and sliding its ray back by one to fill in, we have a natural bijection between $[n]\times\mathbb{N}$, on which $H_{n}$ acts, and $([n]\times\mathbb{N})\setminus\{(i,j)\}$, on which $\operatorname{Stab}(i,j)$ acts, and we see that these actions are the same. (This fact was known to Houghton, see [Hou78].)
@@
To extend these results to Surface Houghton Groups, we shall work out more carefully the embedding $\operatorname{Br}H_{n}\hookrightarrow\mathcal{H}_{n}$ suggested by [ABKL23]. For the braided Houghton group, we take the asymptotically rigid mapping class group of the following surface (see [Fun07] for details). Begin with a $2n$-gon as the center piece, and take a punctured square for the attached pieces (see Figure 4). In [GLU21], it was shown that the braided Houghton group $\operatorname{Br}H_{n}$ is type $F_{n-1}$ but not type $FP_{n}$. This makes it a good candidate for an infinite index subgroup of $\mathcal{H}_{n}$ with $fl(\operatorname{Br}H_{n})=fl(\mathcal{H}_{n})$.
{tikzpicture}
@@
{tikzcd}
By the construction of Ivanov-McCarthy, the map $b$ is an injective homomorphism. The kernel of $a$ is generated by half-twists about the compact boundary components, which are trivial under $b$. Therefore, we obtain an injective homomorphism $\operatorname{Br}H_{n}\rightarrow\mathcal{H}_{n}$. As $\operatorname{Br}H_{n}$ acts trivially on the maximal (i.e. non-puncture) ends, this image lies in $P\mathcal{H}_{n}$.
The image of this homomorphism is a subgroup which fixes (up to isotopy) the set of curves which correspond to the following curve via the canonical maps $\iota_{B}$ of Section 2.1 [Add picture of curve circumscribing the genus of a torus with two boundary components]. All that remains is to see that this subgroup has infinite index in $P\mathcal{H}_{n}$. Consider its intersection with the mapping class group of any suited subsurface: here, it must fix the multicurve consisting of the above curves, and is therefore of infinite index. As infinite index in a subgroup implies infinite index in the full group, we have our result.
It is easy to see that $\operatorname{Br}H_{n}$ is not co-Hopfian: there is an inclusion map from the defining surface of $\operatorname{Br}H_{n}$ to itself, sliding the first puncture in some end out, and with all elements moving things only on one side of the skipped puncture (see Figure 5). As with the Houghton groups, doing such a move in all ends simultaneously yields a homomorphism whose iterated images act trivially on arbitrarily large compact subsurfaces.
{tikzpicture}
Figure 5: Pushing one puncture “off to the side”
These approaches are not immediately available for the surface Houghton group, as there is no inclusion of surfaces which skips over a single genus. In fact, in light of the results of [ALM23], if the pure surface Houghton group were to fail to be co-Hopfian, then it must fail either by a non-twist-preserving homomorphism, or by a homomorphism which is not the restriction of a homomorphism on the level of pure mapping class groups.
All told, we have the following for co-Hopfianness of the asymptotically rigid mapping class group versions of Houghton groups.
@@
Question 5.7.

### v2

@@ 5 Subgroups of maximal finiteness length
We use the notation $fl(G)=n$ to mean that $G$ is a group of type $F_{n}$ but not of type $F_{n+1}$. We know that if $G<H$ is finite index, then $fl(G)=fl(H)$. But what of the converse, that is, when does $fl(G)=fl(H)$ imply that $G$ is finite index in $H$? For Houghton groups, and for the (pure) surface Houghton groups, the answer is positive for sufficiently large subgroups, in particular for coabelian subgroups. We denote by $G^{\prime}$ the commutator subgroup of a group $G$. Recall that the commutator subgroups of the Houghton group $H_{n}$ and the pure surface Houghton group $P\mathcal{H}_{n}$ are the finitely supported and compactly supported elements, respectively. (Really, we could say compactly supported in both cases, using the discrete topology for the $\mathbb{N}$-rays on which the Houghton group acts.) For Theorem 5.1 and Proposition 5.2, let $H$ be either $H_{n}$ or $P\mathcal{H}_{n}$. For $H_{n}$, this is a mild extension of [Zar20, Corollary 2.6].
@@
Let $G<H$ be a subgroup intersecting the commutator subgroup $H^{\prime}$ in a finite index subgroup. If fl(G)=n1fl(G)=n-1, then $G$ is finite index in $H$.
Note that it suffices to show this for groups containing the commutator subgroup. We begin by showing that when $G$ contains the commutator subgroup, the image of $G$ in the abelianization is a maximal rank sublattice, which will then imply finite index.
@@
Let $G<H$ be as above, i.e. fl(G)=n1fl(G)=n-1, and $G$ contains the commutator $H^{\prime}$. Write $F$ for the abelianization map to $\mathbb{Z}^{n-1}$. Then $F(G)$ is finite index in $\mathbb{Z}^{n-1}$.
@@
We now prove Theorem 5.1, using the characterization of the commutator subgroup as the elements of compact support. For $H_{n}$, the ends correspond to the rays in the obvious way.
@@
Suppose first that $G$ contains the commutator subgroup. Consider the map of coset spaces $\Psi:H/G\rightarrow\mathbb{Z}^{n-1}/(\textbf{c}\mathbb{Z}^{n-1})$ given by $hG\mapsto F(h)\textbf{c}\mathbb{Z}^{n-1}$. To see that this is well-defined, suppose $h_{1}G=h_{2}G$: then $h_{1}=h_{2}g$ for some $g\in G$, and $F(h_{1})=F(h_{2})+F(g)$. As $F(g)\in\textbf{c}\mathbb{Z}^{n-1}$, we see that $\Psi(h_{1}G)=\Psi(h_{2}G)$. We now wish to show that $\Psi$ is injective. Suppose that $\Psi(h_{1}G)=\Psi(h_{2}G)$. Then $F(h_{1}h_{2}^{-1})=F(h_{1})-F(h_{2})=F(g)$ for some $g\in G$. So our question becomes: does $F(h_{1}h_{2}^{-1})\in F(G)$ imply $h_{1}h_{2}^{-1}\in G$? The element $h_{1}h_{2}^{-1}g^{-1}$ will have no net translation in any end, and hence is compactly supported, and therefore is in $G$.
As there are clearly subgroups of finite index in both cases (take the pre-image of a finite index subgroup of $\mathbb{Z}^{n-1}$), there is one loose end remaining: the infinite index case. Specifically, do there exist subgroups $G<H_{n}$ (or $G<P\mathcal{H}_{n}$) whose intersection with the commutator subgroup have infinite index, and such that fl(G)=n1fl(G)=n-1. First, we see that $G$ is necessarily infinite index in $H_{n}$ (or $P\mathcal{H}_{n}$), so all that remains is to check whether this case actually occurs. We thank Noel Brady for the proofs of the following lemmas.
@@
Suppose $H:G]=m<\infty$. Write $H=Gh_{1}\cup\cdots\cup Gh_{m}$, as a disjoint union. For each index $i$ such that $K\cap Gh_{i}\neq\emptyset$, there is some $g_{i}\in G$ and $k_{i}\in K$ such that $k_{i}=h_{i}g_{i}$. Then we can write
@@
By the proof of Lemma 5.3, we see that the sequence $([K_{i}:G\cap K_{i}])$ is bounded by $[H:G]=m$. It is not hard to see that it is non-decreasing, so we need only show that it can’t stabilize below $m$. Choosing a decomposition of $H$ into $G$-cosets, say $H=Gh_{1}\cup\cdots\cup Gh_{m}$, we see that there must be some index $j$ such that $K_{j}$ contains all of $\{h_{1},\dots,h_{m}\}$. Therefore, the sequence of indices stabilizes at $m$.
For the converse, consider a decomposition $H=Gh_{1}\cup Gh_{2}\cup\cdots$. Intersecting this with $K_{i}$, we see as before that the index $K_{i}:G\cap K_{i}]$ is the number of indices $j$ such that $Gh_{j}\cap K_{i}$ is non-empty. As every $h_{j}$ must be in some $K_{i}$, we see that if $[H:G]$ is infinite, then the sequence $([K_{i}:G\cap K_{i}])$ goes to infinity. ∎
We return to our question on the existence of infinite index subgroups of Houghton groups with finiteness length n1n-1, We must now split off the ordinary Houghton group from the surface Houghton group. For $H_{n}$, the answer to our existence question is easily yes: there are even copies of $H_{n}$ itself as infinite index subgroups of $H_{n}$! Consider the stabilizer of a single point in $n]\times\mathbb{N}$: ignoring the fixed point, and sliding its ray back by one to fill in, we have a natural bijection between $n]\times\mathbb{N}$, on which $H_{n}$ acts, and $([n]\times\mathbb{N})\setminus\{(i,j)\}$, on which $\operatorname{Stab}(i,j)$ acts, and we see that these actions are the same. (This fact was known to Houghton, see [Hou78].)
@@
To extend the result on the existence of infinite index subgroups of finiteness length n1n-1 to surface Houghton groups, we shall work out more carefully the embedding $\operatorname{Br}H_{n}\hookrightarrow\mathcal{H}_{n}$ suggested by [ABKL23]. For the braided Houghton group, we take the asymptotically rigid mapping class group of the following surface, which we shall call $\dot{\Sigma}$ (see [Fun07] for details). Begin with a $2n$-gon as the center piece, and take a punctured square for the attached pieces (see Figure 4). In [GLU21], it was shown that the braided Houghton group $\operatorname{Br}H_{n}$ is type $F_{n-1}$ but not type $FP_{n}$. This makes it a good candidate for an infinite index subgroup of $\mathcal{H}_{n}$ with $fl(\operatorname{Br}H_{n})=fl(\mathcal{H}_{n})$.
@@
[…]
The failure of $b$ to be an injective homomorphism is precisely the subgroup generated by boundary parallel twists. The kernel of $a$ is generated by twists parallel to the compact boundary components. Therefore, we obtain an injective homomorphism $\operatorname{Map}(\dot{\Sigma})\rightarrow\operatorname{Map}(\Sigma)$. Restricting to asymptotically rigid subgroups yields an injective homomorphism $\operatorname{Br}H_{n}\rightarrow\mathcal{H}_{n}$. As $\operatorname{Br}H_{n}$ acts trivially on the maximal (i.e. non-puncture) ends, this image lies in $P\mathcal{H}_{n}$.
The image of this homomorphism is a subgroup which fixes (up to isotopy) the multicurve defined by the images of the curve $\beta$ in Figure 5 in each piece, via the canonical maps $\iota_{B}$ of Section 2.1.
$\beta$
Figure 5: The central curve $\beta$ in a piece which is fixed by the image of $\operatorname{Br}H_{n}$ in $P\mathcal{H}_{n}$
All that remains is to see that this subgroup has infinite index in $P\mathcal{H}_{n}$. Consider its intersection with the mapping class group of any suited subsurface: here, it must fix the multicurve consisting of the above curves, and is therefore of infinite index. As infinite index in a subgroup implies infinite index in the full group, we have our result.
Between Theorem 5.1, Proposition 5.5, and the above discussion, we have proven the following:
Theorem 5.6.
We finish with further consideration of co-Hopfianness. It is easy to see that $\operatorname{Br}H_{n}$ is not co-Hopfian: there is an inclusion map from the defining surface of $\operatorname{Br}H_{n}$ to itself, sliding the first puncture in some end out, and with all elements moving things only on one side of the skipped puncture (see Figure 6). As with the Houghton groups, doing such a move in all ends simultaneously yields a homomorphism whose iterated images act trivially on arbitrarily large compact subsurfaces. This yields:
@@
Figure 6: Pushing one puncture “off to the side”
These approaches are not immediately available for the surface Houghton group, as there is no inclusion of surfaces which skips over a single genus. In fact, in light of the results of [ALM23], if the pure surface Houghton group were to fail to be co-Hopfian, then it must fail either by a non-twist-preserving homomorphism, or by a homomorphism which is not the restriction of a homomorphism on the level of pure mapping class groups. This leaves us with the following question:
Question 5.8.